An island economy consumes only jellybeans, running shoes, Pepsi, and Vanilla ice compact discs. The base year is 2010. Prices in the base period were $2 per bag of jelly beans, $60 per pair of running shoes, 50 cents per can of Pepsi, and $10 per vanilla ice CD. The amounts purchased were 250 bags of jellybeans, 10 pairs of running shoes, 1000 cans of Pepsi, and 40 Vanilla ice CDs during 2010. In 2013, jellybeans cost $1 per bag, running shoes cost $50 per pair, Pepsi cost 25 cents per can, and Vanilla Ice CDs cost $15 each. Using 2010 as base year, the economy's Consumer Price Index (A Laspeyres Index) for 2013 is_________.(show you work)
Answer #1

Cost of basket in 2010 ($) = 2 x 250 + 60 x 10 + 0.5 x 1000 + 10 x 40 = 500 + 600 + 500 + 400 = 2000

Cost of basket in 2013 ($) = 1 x 250 + 50 x 10 + 0.25 x 1000 + 15 x 40 = 250 + 500 + 250 + 600 = 1600

CPI in 2013 = (Cost of basket in 2013 / Cost of basket in 2010) x 100 = ($1600 / $2000) x 100 = 80